

********************************************************************
UARS WEEKLY STATUS REPORT - 13 AUG 1999
********************************************************************
***********
SOLSTICE
***********
Other than being briefly turned off for the solar eclipse on August 11,
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are completing close-out documentation on
the HEPS Energy Processor trouble; however progress has been hampered
with the departure of Peter Thomas.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The instrument is now off as scheduled.
***********
HRDI
***********
HRDI is operating normally.
***********
WINDII
***********
WINDII was activated yesterday (August 12) and is operating
normally. Data will be acquired until August 15.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
15 days CLA (7.0) PROD (L1-L3) (473R-173C-12F=288L)
45 days GPI (4.05) PROD (CLAES L3B) (95R-90C=5L)
0 days WIN (5.11) PROD (L1-L3) (1540R-1028C-33F=479L)
5 days WIN (5.11A) ENG (L1-L3) (7R-5C=2L)
7 days WIN (5.11A) ENG (LOG) (7R-7C=0L)
The CDHF continued integrating the "UARS" environment on CDHFA/B.
DAAC data created in the month of June 1999 has been previewed for
HALOE, PEM, NMC, and UKMO.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 31 July 1999 (Orbit 43087, GMT Day 212) through 6 August
1999 (Orbit 43191, GMT Day 218).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 218/1634 43187 On
HALOE 218/1732 43188 Off
HALOE 218/2130 43190 On
The beta angle ranged between the angles of 60.2 degrees to 75.1 degrees for
this report period. The maximum beta angle observed for this peak was 75.1
degrees on 31 July 1999 (DOY 212).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.94 to 1.27 degrees C. Additional battery performance data for each
battery on 30 July 1999 and 06 August 1999 are provided in Appendix D.
The clock error ranged from -13.0 to 10.9 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 06
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 04
% HRDI = 03
% HALOE = 03
% SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 04
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/208: 2287.496910
% Transponder B: 99/209: 2287.498901
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/217 show the observatory data loss to be 24 hours, 54
minutes, 46.004 seconds (an increase of 6.272 seconds since the last report
period). This is a 0.0361 percent data loss which equals a 99.9639 percent
data capture for the mission.
The increase of 6.272 seconds occurred over four (4) days:
7 0.224 seconds on DOY 211 due to data hit.
7 1.312 seconds on DOY 212 due to IP line hit.
7 4.224 seconds on DOY 214 due to IP line hit.
7 0.512 seconds on DOY 216 due to data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The current
work is a continuation of the SCS-21 UARS implementation effort done last year
which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to change
battery V/T curves based on loading to mitigate battery risks in sun pointing
safehold.
YEAR 2000 OPERABILITY SUPPORT
The FOT received a set of Y2K planning aids from FDF which will be used to
facilitate UARS Y2K end-to-end testing. Eleven 3-day planning aid files were
received and processed by the FOT. Two of these files did not process
successful on the CMS due to either incorrect data, or invalid data formats.
FDF has been notified of the problems, and are in the process of correcting the
files. (Have not received a status as of yet).
MEETINGS
August 3, 1999 (DOY 215): Y2k meeting to discuss the upcoming end-to-end
testing as well as the MILA test. The backup CMS has to be updated with the
current operations software before any of the testing can be completed, Mike
Hannah has been tasked with action.
August 4, 1999 (DOY 216): Monthly UARS meeting. Combined meeting to present
the annual history of UARS and upcoming events. FOT was assigned five action
items.
1) Understand the manetometer/FSS safehold recovery procedure including
recovery procedures from safehold.
2) Understand limitations and operations of crippled solar array drive.
3) Performance trending of NBTR-B
4) Generate annular Leonid meteor shower plan.
5) Investigate UARS end of life disposal plan.
August 5, 1999 (DOY 217): UARS Control Center Meeting was held to discuss the
DOC hardware. FOT was assigned two Action items:
1) Write a report on the impact of the removal of the RUPS, due August 19th.
2) Figure out how many of the machines in the DOC (APUs, TACUs, etc..._ are
really needed to fly our spacecraft.
Post meeting, Dimitrios gave Charlie Coleman the requirements for what will be
needed to complete the upcoming simulator testing.
August 6, 1999 (DOY 218): FOT met to discuss final preparations for partial
solar eclipse.
FUTURE SIGNIFICANT EVENTS
August 20, 1999 (DOY 232): The next UARS Yaw Around Maneuver (forward-to
reverse) is scheduled.
Aug 11, 1999 (DOY 223): Preparations for a partial eclipse of the Sun is
planned. All instruments (except SUSIM) will be commanded off two orbits prior
to the eclipse (0900 Zulu) and then power on two orbits following the event
(1700 Zulu). Final details are available.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 6 August 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
43115 213/2128 99-266 HRDI Telescope Safehold - hi temp
43129 214/1938 99-267 SUSIM Turned Detector off
43135 215/0533 99-268 OBC Clk Adj fr 41.2 to 20.6 msec/day
43140 215/1215 99-269 HRDI HRDI Telescope Safe - Hi Bar Temp
43141 215/1358 99-270 HRDI HRDI Reconfigure for Fulltime
43163 217/0202 99-271 OBC Clk Adj fr 20.6 to 41.2 msec/day
43180 218/0525 99-272 OBC Clk Adj 41.2 to 20.6 msec/day
43184 218/1044 99-273 OBC Clk Adj 20.6 to 0.00 msec/day
43187 218/1634 99-274 HALOE Power ON
43188 218/1732 99-275 HALOE Power OFF
43190 218/2130 99-276 HALOE Power ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 6 August 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSBARTMP Y-LO HRDI BARREL TEMP 43108-13
HRPTSBARTMP R-LO HRDI BARREL TEMP 43114-15
HRPTSBARTMP Y-LO HRDI BARREL TEMP 43135-39,42
HRPTSBARTMP R-LO HRDI BARREL TEMP 43140-41
HRIFESURVTMP Y-HI HRDI SURVIVAL TEMP 43149-51
HRRADTEMP Y-HI HRDI RADIATOR TEMP 43151
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 6 August 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
43091 212/06:03 (31 July 99) from 20.6 to 41.2 msec/day
43135 215/05:33 (3 August 99) from 41.2 to 20.6 msec/day
43163 217/02:08 (3 August 99) from 20.6 to 41.2 msec/day
43180 218/05:25 (3 August 99) from 41.2 to 20.6 msec/day
43184 218/10:44 (3 August 99) from 20.6 to 0.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 6 August 1999
30 July 1999 - GMT Day 211 Beta = 74.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ --------- ----------- ----- ------- ------
1 -112.0/-134.4 -2.42/-3.68 +0.0/+0.0 16.5 0.000 0.0
2 +22.4/-5.6 +2.60/+1.66 +16.4/-13.6 27.5 1.036 13.4
3 +22.4/-22.4 +1.97/+1.03 +16.0/-13.6 27.5 1.035 13.1
6 August 1999 - GMT Day 218 Beta = 60.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ----- ------- ------
1 -112.0/-134.4 -3.05/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +22.4/-11.2 +4.51/+2.60 +23.2/-15.2 26.4 1.025 17.2
3 +33.6/-16.8 +3.23/+1.66 +23.2/-14.4 26.2 1.020 16.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 6 August 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT 43139 215/1026 14170 Generic Late Acquisition
99-039 FOT 43175 217/2052 14170 Generic Late Acquisition
99-040 FOT 43187 218/1634 21201 NASCOM problem at WSGT
99-039 FOT 43189 218/1949 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT see list July 99 14170 Generic Late Acq.
99-040 FOT 43187 218/1634 21201 NASCOM problem at WSGT
SEVNTFW99/32